OpenStudy (anonymous):

If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4).

OpenStudy (anonymous):

-4x-4

OpenStudy (campbell_st):

well f - g is just f(x) - g(x) which is 2x - 2 - (3x -1) = -x -1 then substitute x = 4